# MemoFurther Readings

### v1

@@ Results
Table 5. Keyword statistics in coordinated and organic content related to QAnon conspiracies on Telegram.
$\mathbb{X}$/Twitter: On Twitter, we identified 221 coordinated users. Figure 2(b) displays the top 10 topics shared by these coordinated accounts, all of which pertain to social, economic, or political issues. The most prevalent topic — and one that diverges significantly from those shared by the organic group — focuses on the Independent candidate Robert F. Kennedy, with a subset of coordinated accounts promoting him as a serious contender in the U.S. Presidential Election. Coordinated users also concentrate on the Democratic Primary race in Westchester County (labeled as “bowman-lattimer” in the figure), where the most representative tweets endorse candidate Lattimer. This discussion appears to be framed around the anti-Israel stance of the incumbent, Bowman. The remaining topics revolve around polarizing issues such as COVID-19 vaccines, U.S. border security, and various scandals, including the Epstein files and the second son of Joe Biden (Hunter Biden). Also, these users discuss perceived opposition within the Black community to President Biden, critiques of the student loan forgiveness policy, and concerns over the Biden administration’s sale of oil reserves. No conspiracy theories were detected among the coordinated users, and only a minimal amount of QAnon-related content was found within the organic group. The most frequent keyword, “qanon”, appeared just five times.
Facebook: We identify 16 coordinated users in the Facebook discourse. As shown in Figure 2(c), most topics promoted by these users are left-leaning. A manual inspection reveals that all 16 accounts are associated with The Young Turks network [^16] https://en.wikipedia.org/wiki/The_Young_Turks, which MBFC classifies as left-leaning [^17] https://mediabiasfactcheck.com/the-young-turks/. Apart from a news item quoting a GOP representative accusing Biden of using “supplements” (i.e., “jacked-up Biden”), we found no trace of misleading, conspiratorial, or inflammatory content.
(a) Telegram
(b) Twitter
(c) Facebook
Figure 3. Empirical Cumulative Distribution of the generated engagement by coordinated and organic users.
Generated Engagement and AI-Generated Content
@@
Figure 4. Distribution of AIGC shared by coordinated and organic users across the three different platforms.
For each platform, we begin by merging the sets of coordinated actors identified through co-URL sharing (web domain promotion) and content similarity patterns (content amplification). We observe that the intersection between these two patterns is modest, with overlap rates of 1.3%, 0.5%, and 0% for Telegram, Twitter, and Facebook, respectively. This is in line with previous findings showing that different strategies are used to push diverse agendas through distinct coordinated networks (Ng and Carley, 2022; Luceri et al., 2024a). Aggregating coordinated actors from distinct CoIA networks, we obtain 233 coordinated actors on Telegram, 764 on Twitter, and 25 on Facebook.
Figure 3 shows the distribution of engagement generated by coordinated and organic users across all platforms. In each case, coordinated users tend to generate less engagement than their organic counterparts. This is expected, as the organic group includes major media outlets and influencers that significantly boost overall engagement. However, it is noteworthy that the engagement from coordinated actors is substantial, and particularly on Twitter and Facebook, the engagement distributions of coordinated and organic users are quite similar.
We also assess the prevalence of AIGC that was shared across the three platforms by both coordinated and organic users. These results are shown in Figure 4. Facebook is the social media platform where most AIGC is produced by both coordinated and organic accounts. While the AIGC prevalence is limited, we observe a striking difference in the activity between coordinated and organic users on Telegram and Facebook, whereas it is quite balanced on Twitter. Interestingly, AI-generated content is predominantly diffused by coordinated actors on Telegram, while an opposite trend is observed on Facebook.
Conclusions
@@
Our unsupervised, network-based methodology allows us to identify a range of influence campaigns aimed at directing traffic toward specific narratives and domains. These campaigns frequently promote content that is partisan, low-credibility, and conspiratorial in nature. The model not only detects domestic actors but also reveals the systematic promotion of foreign-affiliated media, particularly Russian state-sponsored outlets, across platforms like Telegram and $\mathbb{X}$. This evidence points to a coordinated effort to amplify certain messages and domains across distinct platforms, particularly in the context of the 2024 U.S. Election, underscoring the necessity of a unified, cross-platform approach to tackle such influence operations (Gatta et al., 2023).
Limitations
While our study demonstrates substantial strengths, a few considerations remain. First, although the model leverages data from Telegram, Facebook, and $\mathbb{X}$, the specific timeframe under observation and platform-specific mechanisms may influence the broader applicability of our model and the generalizability of our findings to other scenarios and social media networks. Second, combining similarity networks with differing similarity distributions across platforms may introduce biases in the overall detection of coordination, potentially favoring certain platforms. Finally, this work does not examine multimedia content, thereby overlooking coordination across different data modalities and the potential use of generative AI techniques for producing synthetic media.
Future work should address these limitations by incorporating sophisticated statistical models to assess the significance of coordination patterns, thereby providing a more robust and statistically sound methodology to ensure that identified coordination signals are not merely artifacts of underlying data distributions. Expanding the framework to encompass additional platforms, data modalities, and a broader range of behavioral signals would further strengthen its capability to detect diverse forms of coordinated inauthentic activity.
About the Team
[…]
Appendix
1. Dataset Collection
@@
(c) Telegram posts (d) Twitter characters
@@
(e) Twitter posts (f) Telegram characters
Table 6. Distributions of posts and characters across platforms.
2. AI Detection Classifier Validation
We utilized the AI-generated text detection classifier from (Dmonte et al., 2024), originally trained to identify AI-generated content in tweets. To extend the applicability of this classifier to other platforms, we constructed an external validation set using diverse, older datasets. Specifically, we built upon four datasets: Twitter-2010181818https://archive.org/details/twitter_cikm_2010, Facebook [^19] https://www.kaggle.com/datasets/sheenabatra/facebook-data, and Telegram.
From each of these datasets, we sampled 500 texts ranging in length from 125 to 1000 characters. Given the release years of these datasets, we assume that all texts are non-AI generated. To create the AI-generated class, we used GPT-4o and Llama 3.1 3B Instruct, prompting them to generate new texts with similar topics and lengths as the original non-AI texts. We performed this generation for each non-AI text and sampled 250 AI-generated texts from GPT and 250 from Llama.
@@ 2. AI Detection Classifier Validation
To further enhance the generalizability of the validation dataset, we incorporated non-AI generated content from additional sources: Reddit comments [^20] https://zissou.infosci.cornell.edu/convokit/datasets/reddit-coarse-discourse-corpus/, IMDB reviews [^21] https://ai.stanford.edu/~amaas/data/sentiment/, movies corpus [^22] https://zissou.infosci.cornell.edu/convokit/datasets/movie-corpus/, and Yelp 2013232323https://www.yelp.com/dataset/download. AI-generated content for these datasets was sourced from tweetHunter [^24] https://tweethunter.io/ and GPT, which was prompted to generate texts on specific topics such as American football, climate change, the Soccer World Cup, the Gaza conflict, U.S. politics, and vaccines.
The validation results are displayed in Table 7.
@@
Table 7. Precision and Recall for each platform
(a) Facebook.
(b) Twitter.
(c) Telegram.
Figure 5. Thresholds grid search for filtering similarity graph: x-axis edge similarity quantile, y-axis node centrality quantile. z-axis is the minimum graph density across all connected components of the filtered co-url similarity graph. The green square corresponds to the selected thresholds.
3. Grid Search for Similarity and Centrality Thresholds
To identify coordinated accounts in the co-URL similarity network, we filter edges and nodes based on cosine similarity and eigenvector centrality. Two nodes are connected if they co-share URLs, and the strength of these connections is represented by the cosine similarity between their TF-IDF vectors across the space of unique URLs.
We employ two filtering techniques commonly found in the literature: edge filtering, based on cosine similarity, and node filtering, based on eigenvector centrality. The thresholds for these filters are determined by the percentiles of the distributions of edge similarity and node centrality.
Assuming that coordinated accounts manifest as dense components in the similarity graph, we use the density of each connected component as a key indicator of coordination. To ensure robustness, we adopt a conservative approach, using the minimum density across all connected components after filtering as a comprehensive quality measure of the graph.
Figure 5 presents the results of our grid search across these parameters. The x-axis represents the percentile of edge similarity, while the y-axis corresponds to the percentile of eigenvector centrality. The z-axis indicates the minimum density of the similarity graph after filtering based on the selected x-y percentiles.
For each platform, we identified the optimal thresholds by observing sharp changes in minimum density and high overall density values. These threshold combinations define the coordinated accounts on each platform. Specifically:
@@ 3. Grid Search for Similarity and Centrality Thresholds
  • For Twitter, we chose (85%, 99%), consistent with previous studies, leading to an increase from 0.74 to 0.99.
  • For Reddit, we selected (85%, 80%), where the minimum density shifted from 0.81 and 0.89 to 0.96.
@@
Figure 6. Thresholds grid search for filtering the Cross-platform similarity graph: x-axis edge similarity quantile, y-axis node centrality quantile. z-axis is the minimum graph density across all connected components of the filtered co-url similarity graph. The green square corresponds to the selected thresholds.
4. Characterization of coordinated users
@@
Table 8. Current bios of top-5 coordinated channels in Telegram sorted by degree.
@@
Table 9. Current bios of top-5 coordinated accounts in Twitter sorted by degree.
@@
Table 10. Current bios of top-5 coordinated pages in Facebook sorted by degree.
@@
Table 11. Current bios of top-5 cross-platform coordinated accounts sorted by degree.
@@
"US Secretary of State Antony Blinken Walking the Congressional Office Building Halls Surrounded by Police ’War criminal’, ’genocide secretary’ Protestors reminding him of his complicity in crimes against humanity. RealWorldNewsChannel RealWorldNewsChat"
"Its sickening, disgusting money laundering MF’s. In a meantime the people of United States are suffering! Can’t afford basic necessities!!"
"I like how zoinists shout poor me when they get caught out for there evil crimes Normal Jews hate zoinist Jews People of the world don’t hate Jews People of the world hate zoinists Jews n zoinists in general n what they stand for They think there the chosen ones and that gives them the rite to kill n genecide inercent people Zoinisium is hated by all races of the world And has no place for peace in humanity "
"Ukrainians in Galway are being advised to vote for two Nigerians and a Labour candidate to best serve their interests. ’Under no circumstances vote for radicals - Irish Freedom Party, Independent Ireland, The Irish People and all others who have the slogans ’Ireland for the Irish’. #ForeignInterference "
"Former CNN Anchor Chris Cuomo Admits to Suffering from a COVID Vaccine Injury ICYMI: There’s been a major shift in the official narrative. Follow Vigilant_News "
Table 12. Top-5 messages by total engagement for coordinated channels in Telegram
@@
"“WE THE PEOPLE” don’t play by your dictator rules. Just answer the damn question, crybaby. White House correspondents fire back after Biden snaps at reporter for refusing to ’play by the rules’"
"If Biden wins there will no more arguing because the will not be anything left to argue about. He’s handing the keys over to illegals and the new world order. In the end, the joke will be on the democrats. The WHO/WEF hates them too and knows they’re imbeciles. Everyone knows"
"’Shameful’: GOP lawmaker shreds ’AWOL’ Biden for throwing Jews ’under the bus’ amid anti-Israel protests If this president, so-called president doesn’t personify evil destruction division of this country, I don’t know what or who would !"
"Biden DHS docs suggested Trump supporters, military and religious people are likely violent terror threats. HaHaHa! I guess we also have Santa Claus/Easter bunny over here in training. Give me a break, we are trying to save this country from the bad guys!"
"DeSantis spox dunks on NYT ’fact-check’ on terrorists entering southern border: ’Awaiting your correction’ They’ve been entering for last 3 years. Biden admin has no damn idea who is coming in. Even ones they think they know are using others’ identity"
Table 13. Top-5 posts by total engagement for coordinated accounts in Twitter
@@
"Youth Vote Shifts Toward Trump in 2024 Election https://gorightnews.com/youth-vote-shifts-toward-trump-in-2024-election/ Maybe the kids are alright… #GoRightNews Recent polls indicate a surprising trend: young voters are warming to Donald Trump in the 2024 presidential election. The question arises: Are these voters aligning with Trump’s policies, or is President Biden driving them away? The answer might lie in a combination of both factors. Polling Data Reveals Shift According to the latest New York Times poll, young voters aged 18 to 29 favor Biden by a slim margin of two points, 47% to 45%. A Quinnipiac poll shows Trump leading Biden among voters aged 18 to 34, 48% to 47%. This is a stark contrast to the 2020 presidential election, where Joe Biden secured the youth vote by a significant 24%. The last time a Republican won this demographic was in 1988. Biden’s Struggling Message Aidan Kohn-Murphy, founder of Gen Z for Change—a group that supported Biden in 2020—stated in the Washington Post, “Biden is out of step with young people on a number of key issues." Key issues where Biden seems to be losing support include: The War in Gaza: A majority of young voters, 51%, support the Palestinians, while only 15% support Israel. TikTok Ban: Biden’s support for a TikTok ban is perceived as an attack on free speech. Consequently, 67% of Gen Z voters say this makes them less likely to vote for him. Economic Challenges: High inflation and interest rates have made essential costs like food and housing unaffordable for young people entering the workforce or trying to purchase their first home. Trump’s Resonating Message The 45th president’s message appears to be resonating with young voters for several reasons: Gaza Conflict: While Trump supports Israel, he promises to bring a peaceful end to the conflict, citing his administration’s four years of peace as evidence of his capability. TikTok Engagement: Trump recently joined TikTok, emphasizing that Biden wants to shut the platform down, thus appealing to younger users. Economic Performance: When asked about Trump’s handling of the economy, 65% of young voters approved, compared to just 33% for Biden. Social Media Dynamics According to CredoIQ, a social media analytics firm, nearly 25% of the top left-leaning content creators on TikTok have posted anti-Biden content in the first four months of 2024, garnering over 100 million views. This content is often created by young, non-white liberals who share the belief that the U.S. Government, and specifically Joe Biden, aims to restrict free speech and information flow. Trump’s Adaptation to New Media In 2016, Trump broke political norms with an ’anyplace, anytime’ approach, dominating cable TV. For the 2024 campaign, he has adapted this strategy to new media, appearing on podcasts, YouTube shows, and attending live events such as UFC and Formula 1. He has also made campaign stops at local venues, including bodegas, firehouses, and even in the South Bronx. The shift in the youth vote suggests a significant realignment in political affiliations, one that underscores the importance of addressing the issues most pertinent to young Americans. This analysis highlights the potential impact of these changing dynamics on the upcoming presidential election. [Source: Washington Post, Axios] https://archive.is/kk3HU https://www.axios.com/2024/06/13/trump-election-young-voters-polling #GoRightNews Shared by Peter Boykin - American Political Commentator / Citizen Journalist / Activist / Constitutionalist for Liberty Web: https://PeterBoykin.com Kick: http://Kick.com/PeterBoykin YouTube: https://youtube.com/PeterBoykinForAmerica Twitter: https://twitter.com/GoRightNews Telegram: http://t.me/realpeterboykin Rumble: http://Rumble.com/GoRightNews Like the Content? Please Support! - Go Right News: Stripe: https://gorightnews.com/donations/support-gorightnews/ Cash App: http://cash.app/\$PeterBoykin1"

### v2

@@ 5.2. Detecting CoIA for Content Amplification (RQ2)
Table 5. QAnon-related keywords promoted by coordinated and organic channels on Telegram. Note that coordinated channels account for only 0.36% of all Telegram channels.
$\mathbb{X}$/Twitter: On $\mathbb{X}$/Twitter, we identified 221 coordinated users. Figure 2(b) displays the top 10 topics shared by these coordinated accounts, all of which pertain to social, economic, or political issues. The most prevalent topic—and one that differs significantly from those shared by the organic group—focuses on Independent candidate Robert F. Kennedy (RFK), with a subset of coordinated accounts actively promoting his candidacy in the U.S. Presidential Election. Coordinated users also concentrate on the Democratic Primary race in Westchester County (labeled as “bowman-lattimer” in Figure 2(b)), where the most representative tweets endorse candidate Lattimer. This discussion appears to be framed around the anti-Israel stance of the incumbent, Rep. Jamaal Bowman. The remaining topics revolve around polarizing issues such as COVID-19 vaccines, U.S. border security, and various scandals, including the Epstein files and Hunter Biden. Also, these users discuss perceived opposition within the Black community to President Biden, critiques of the student loan forgiveness policy, and concerns over the Biden administration’s sale of oil reserves. It should be noted that the most representative tweets for each topic (as identified by BERTopic) were manually reviewed by three authors of this paper, who found no evidence of conspiratorial content (see Table E.11 in the Appendix). The most frequent QAnon-related keyword, “qanon,” appeared only five times.
Facebook: We identify 16 coordinated users in the Facebook discourse. As shown in Figure 2(c), most topics promoted by these users are left-leaning. A manual inspection reveals that all 16 accounts are associated with The Young Turks network171717en.wikipedia.org/wiki/The_Young_Turks, which MBFC classifies as left-leaning181818mediabiasfactcheck.com/the-young-turks/. Apart from a news item quoting a GOP representative accusing Biden of using “supplements” (i.e., “jacked-up Biden”), we found no trace of misleading, conspiratorial, or inflammatory content.
Refer to caption Refer to caption Refer to caption
Figure 3. Empirical Cumulative Distribution of total engagement, defined as the sum of user interactions across platforms, including reactions, comments, shares, replies, and views—generated by coordinated and organic users.
5.3. Generated Engagement and AIGC (RQ3)
@@
Figure 4. AIGC distribution across the three platforms, highlighting the prevalence of AI-generated posts among coordinated and organic users.
To maintain a consistent definition of engagement across platforms, we define it as the sum of likes, replies, quotes, and retweets for $\mathbb{X}$/Twitter, likes, comments, and shares for Facebook, and views, forwards, and replies for Telegram.
For each platform, we begin by merging the sets of coordinated actors identified through co-URL sharing (web domain promotion) and content similarity patterns (content amplification). We observe that the overlap between these two sets of coordinated users is modest, with overlap rates of 1.3%, 0.5%, and 0% for Telegram, Twitter, and Facebook, respectively. This aligns with previous findings showing that different coordination strategies are employed to push diverse agendas through distinct coordinated networks (Ng and Carley, 2022; Luceri et al., 2024b). This observation supports our decision to apply two complementary detection strategies. Aggregating coordinated actors from distinct CoIA networks, we identify 233 coordinated actors on Telegram, 764 on $\mathbb{X}$, and 25 on Facebook. The proportion of accounts engaged in CoIA is 1.49% on Telegram, 0.43% on $\mathbb{X}$, and 0.41% on Facebook, consistent with prior research that employed conservative thresholds to detect coordinated activity (Pacheco et al., 2021; Luceri et al., 2024b).
Table 3 shows the distribution of engagement generated by coordinated and organic users across all platforms. In each case, coordinated users tend to generate less engagement than their organic counterparts. This is expected, as the organic group includes major media outlets and influencers that significantly boost overall engagement. However, it is noteworthy that the engagement from coordinated actors is substantial, and particularly on $\mathbb{X}$ and Facebook, the engagement distributions of coordinated and organic users are quite similar.
We also assess the prevalence of AIGC that was shared across the three platforms by both coordinated and organic users. These results are shown in Figure 4. Facebook is the social media platform where most AIGC is produced by both coordinated and organic accounts. While the AIGC prevalence is limited, we observe a striking difference in the activity between coordinated and organic users on Telegram and Facebook, whereas it is quite balanced on $\mathbb{X}$. Interestingly, AI-generated content is predominantly diffused by coordinated actors on Telegram, while an opposite trend is observed on Facebook.
6. Conclusions
@@
Our unsupervised, network-based methodology allows us to identify a range of influence campaigns aimed at directing traffic toward specific narratives and domains. These campaigns frequently promote content that is partisan, low-credibility, and conspiratorial in nature. The model not only detects domestic actors but also reveals the systematic promotion of foreign-affiliated media, particularly Russian state-sponsored outlets, across platforms like Telegram and $\mathbb{X}$. This evidence points to a coordinated effort to amplify certain messages and domains across distinct platforms underscoring the necessity of a unified, cross-platform approach to tackle such influence operations (Gatta et al., 2023).
Limitations. While our study demonstrates substantial strengths, a few considerations remain. First, although the model leverages data from Telegram, Facebook, and $\mathbb{X}$, the specific timeframe under observation and platform-specific mechanisms may influence the broader applicability of our model and the generalizability of our findings to other scenarios and social media networks. Second, combining similarity networks with differing similarity distributions across platforms may introduce biases in the overall detection of coordination, potentially favoring certain platforms. Finally, this work does not examine multimedia content, thereby overlooking coordination across different data modalities and the potential use of generative AI techniques for producing synthetic media.
Future work should address these limitations by incorporating sophisticated statistical models to assess the significance of coordination patterns, thereby providing a more robust and statistically sound methodology to ensure that identified coordination signals are not merely artifacts of underlying data distributions. Expanding the framework to encompass additional platforms, data modalities, and a broader range of behavioral signals would further strengthen its capability to detect diverse forms of coordination.
Acknowledgments
This work was supported in part by DARPA (#HR001121C0169). MM acknowledges partial support by the SERICS project (PE00000014) under the NRRP MUR program funded by the EU - NGEU.
Appendix A Dataset Collection
@@
(c) Telegram posts (d) $\mathbb{X}$ characters
@@
(e) $\mathbb{X}$ posts (f) Telegram characters
Table A.1. Distributions of posts and characters across platforms.
Appendix B AI Detection Classifier Validation
We utilized the AI-generated text detection classifier from (Dmonte et al., 2024), originally trained to identify AI-generated content in tweets. To extend the applicability of this classifier to other platforms, we constructed an external validation set using diverse, older datasets. Specifically, we built upon four datasets: Twitter-2010191919https://archive.org/details/twitter_cikm_2010, Facebook [^20] https://www.kaggle.com/datasets/sheenabatra/facebook-data, and Telegram.
From each of these datasets, we sampled 500 texts ranging in length from 125 to 1000 characters. Given the release years of these datasets, we assume that all texts are non-AI generated. To create the AI-generated class, we used GPT-4o [^21] https://openai.com/index/hello-gpt-4o/ and Llama 3.1 3B Instruct [^22] https://huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct, prompting them to generate new texts with similar topics and lengths as the original non-AI texts. We performed this generation for each non-AI text and sampled 250 AI-generated texts from GPT and 250 from Llama.
@@ Appendix B AI Detection Classifier Validation
To further enhance the generalizability of the validation dataset, we incorporated non-AI generated content from additional sources: Reddit comments [^23] https://zissou.infosci.cornell.edu/convokit/datasets/reddit-coarse-discourse-corpus/, IMDB reviews [^24] https://ai.stanford.edu/~amaas/data/sentiment/, movies corpus [^25] https://zissou.infosci.cornell.edu/convokit/datasets/movie-corpus/, and Yelp 2013262626https://www.yelp.com/dataset/download. AI-generated content for these datasets was sourced from tweetHunter [^27] https://tweethunter.io/ and GPT, which was prompted to generate texts on specific topics such as American football, climate change, the Soccer World Cup, the Gaza conflict, U.S. politics, and vaccines.
The validation results are displayed in Table B.2.
@@
Table B.2. Precision and Recall for each platform
Appendix C Grid Search for Similarity and Centrality Thresholds
To identify coordinated accounts in the co-URL similarity network, we filter edges and nodes based on cosine similarity and eigenvector centrality. Two nodes are connected if they co-share URLs, and the strength of these connections is represented by the cosine similarity between their TF-IDF vectors across the space of all unique URLs. We employ two filtering techniques commonly found in the literature: edge filtering, based on cosine similarity, and node filtering, based on eigenvector centrality. The thresholds for these filters are determined by the percentiles of the distributions of edge similarity and node centrality. Assuming that coordinated accounts manifest as dense components in the similarity graph, we use the density of each connected component as a key indicator of coordination. To ensure robustness, we adopt a conservative approach, using the minimum density across all connected components after filtering as a comprehensive quality measure of the graph. Figure C.1 presents the results of our grid search across these parameters. The x-axis represents the percentile of edge similarity, while the y-axis corresponds to the percentile of eigenvector centrality. The z-axis indicates the minimum density of the similarity graph after filtering based on the selected x-y percentiles. For each platform, we identified the optimal thresholds by observing sharp changes in minimum density and high overall density values. These threshold combinations define the coordinated accounts on each platform. Specifically:
@@ Appendix C Grid Search for Similarity and Centrality Thresholds
  • For $\mathbb{X}$, we chose (85%, 99%), consistent with previous studies, leading to an increase from 0.74 to 0.99.
@@
Performance Evaluation. To validate our methodology, we used the supervised Information Operations (IO) dataset from (Seckin et al., 2024), which includes labeled users identified as coordinated inauthentic actors across 26 IO campaigns, along with a control group of organic users. Our method achieves high precision (above 0.9) but low recall (around 0.1), aligning with the framework’s design and the objectives of our analysis.
(a) Facebook.
(b) $\mathbb{X}$.
(c) Telegram.
Figure C.1. Thresholds grid search for filtering similarity graph: x-axis edge similarity quantile, y-axis node centrality quantile. z-axis is the minimum graph density across all connected components of the filtered co-url similarity graph. The green square corresponds to the selected thresholds.
Figure C.2. Thresholds grid search for filtering the Cross-platform similarity graph: x-axis edge similarity quantile, y-axis node centrality quantile. z-axis is the minimum graph density across all connected components of the filtered co-url similarity graph. The green square corresponds to the selected thresholds.
Appendix D Characterization of coordinated users
@@
Table D.3. Current bios of top-5 coordinated channels in Telegram sorted by degree.
@@
Table D.4. Current bios of top-5 coordinated accounts in $\mathbb{X}$ sorted by degree.
@@
Table D.5. Current bios of top-5 coordinated pages in Facebook sorted by degree.
@@
Table D.6. Current bios of top-5 cross-platform coordinated accounts sorted by degree.
@@
”US Secretary of State Antony Blinken Walking the Congressional Office Building Halls Surrounded by Police ’War criminal’, ’genocide secretary’ Protestors reminding him of his complicity in crimes against humanity. RealWorldNewsChannel RealWorldNewsChat”
”Its sickening, disgusting money laundering MF’s. In a meantime the people of United States are suffering! Can’t afford basic necessities!!”
”I like how zoinists shout poor me when they get caught out for there evil crimes Normal Jews hate zoinist Jews People of the world don’t hate Jews People of the world hate zoinists Jews n zoinists in general n what they stand for They think there the chosen ones and that gives them the rite to kill n genecide inercent people Zoinisium is hated by all races of the world And has no place for peace in humanity ”
”Ukrainians in Galway are being advised to vote for two Nigerians and a Labour candidate to best serve their interests. ’Under no circumstances vote for radicals - Irish Freedom Party, Independent Ireland, The Irish People and all others who have the slogans ’Ireland for the Irish’. #ForeignInterference ”
”Former CNN Anchor Chris Cuomo Admits to Suffering from a COVID Vaccine Injury ICYMI: There’s been a major shift in the official narrative. Follow Vigilant_News ”
Table D.7. Top-5 messages by total engagement for coordinated channels in Telegram
@@
”“WE THE PEOPLE” don’t play by your dictator rules. Just answer the damn question, crybaby. White House correspondents fire back after Biden snaps at reporter for refusing to ’play by the rules’”
”If Biden wins there will no more arguing because the will not be anything left to argue about. He’s handing the keys over to illegals and the new world order. In the end, the joke will be on the democrats. The WHO/WEF hates them too and knows they’re imbeciles. Everyone knows”
”’Shameful’: GOP lawmaker shreds ’AWOL’ Biden for throwing Jews ’under the bus’ amid anti-Israel protests If this president, so-called president doesn’t personify evil destruction division of this country, I don’t know what or who would !”
”Biden DHS docs suggested Trump supporters, military and religious people are likely violent terror threats. HaHaHa! I guess we also have Santa Claus/Easter bunny over here in training. Give me a break, we are trying to save this country from the bad guys!”
”DeSantis spox dunks on NYT ’fact-check’ on terrorists entering southern border: ’Awaiting your correction’ They’ve been entering for last 3 years. Biden admin has no damn idea who is coming in. Even ones they think they know are using others’ identity”
Table D.8. Top-5 posts by total engagement for coordinated accounts in $\mathbb{X}$
@@
”Youth Vote Shifts Toward Trump in 2024 Election https://gorightnews.com/youth-vote-shifts-toward-trump-in-2024-election/ Maybe the kids are alright… #GoRightNews Recent polls indicate a surprising trend: young voters are warming to Donald Trump in the 2024 presidential election. The question arises: Are these voters aligning with Trump’s policies, or is President Biden driving them away? The answer might lie in a combination of both factors. Polling Data Reveals Shift According to the latest New York Times poll, young voters aged 18 to 29 favor Biden by a slim margin of two points, 47% to 45%. A Quinnipiac poll shows Trump leading Biden among voters aged 18 to 34, 48% to 47%. This is a stark contrast to the 2020 presidential election, where Joe Biden secured the youth vote by a significant 24%. The last time a Republican won this demographic was in 1988. Biden’s Struggling Message Aidan Kohn-Murphy, founder of Gen Z for Change—a group that supported Biden in 2020—stated in the Washington Post, “Biden is out of step with young people on a number of key issues.” Key issues where Biden seems to be losing support include: The War in Gaza: A majority of young voters, 51%, support the Palestinians, while only 15% support Israel. TikTok Ban: Biden’s support for a TikTok ban is perceived as an attack on free speech. Consequently, 67% of Gen Z voters say this makes them less likely to vote for him. Economic Challenges: High inflation and interest rates have made essential costs like food and housing unaffordable for young people entering the workforce or trying to purchase their first home. Trump’s Resonating Message The 45th president’s message appears to be resonating with young voters for several reasons: Gaza Conflict: While Trump supports Israel, he promises to bring a peaceful end to the conflict, citing his administration’s four years of peace as evidence of his capability. TikTok Engagement: Trump recently joined TikTok, emphasizing that Biden wants to shut the platform down, thus appealing to younger users. Economic Performance: When asked about Trump’s handling of the economy, 65% of young voters approved, compared to just 33% for Biden. Social Media Dynamics According to CredoIQ, a social media analytics firm, nearly 25% of the top left-leaning content creators on TikTok have posted anti-Biden content in the first four months of 2024, garnering over 100 million views. This content is often created by young, non-white liberals who share the belief that the U.S. Government, and specifically Joe Biden, aims to restrict free speech and information flow. Trump’s Adaptation to New Media In 2016, Trump broke political norms with an ’anyplace, anytime’ approach, dominating cable TV. For the 2024 campaign, he has adapted this strategy to new media, appearing on podcasts, YouTube shows, and attending live events such as UFC and Formula 1. He has also made campaign stops at local venues, including bodegas, firehouses, and even in the South Bronx. The shift in the youth vote suggests a significant realignment in political affiliations, one that underscores the importance of addressing the issues most pertinent to young Americans. This analysis highlights the potential impact of these changing dynamics on the upcoming presidential election. [Source: Washington Post, Axios] https://archive.is/kk3HU https://www.axios.com/2024/06/13/trump-election-young-voters-polling #GoRightNews Shared by Peter Boykin - American Political Commentator / Citizen Journalist / Activist / Constitutionalist for Liberty Web: https://PeterBoykin.com Kick: http://Kick.com/PeterBoykin YouTube: https://youtube.com/PeterBoykinForAmerica Twitter: https://twitter.com/GoRightNews Telegram: http://t.me/realpeterboykin Rumble: http://Rumble.com/GoRightNews Like the Content? Please Support! - Go Right News: Stripe: https://gorightnews.com/donations/support-gorightnews/ Cash App: http://cash.app/\$PeterBoykin1”